OpenStudy (anonymous):

What is the solution to the compound inequality below? 7x+3/4≥13 or 5/2x-1/3≥-11/2

OpenStudy (anonymous):

I have a few answers to choose from but I am having trouble with all the fractions!

OpenStudy (imstuck):

\[7x+\frac{ 3 }{4 }\ge 13 \]or\[\frac{ 5 }{ 2 }x-\frac{ 1 }{ 3 }\ge -\frac{ 11 }{ 2 }\]Is it that?

OpenStudy (anonymous):

these are my choices

OpenStudy (anonymous):

thanks so much for helping me though!

OpenStudy (imstuck):

It's b. I can show you how to do them, ok?

OpenStudy (imstuck):

Pay attention.

OpenStudy (imstuck):